OpenStudy (anonymous):

If P (A) = 0.36, what is P (Ā)?

OpenStudy (anonymous):

The probability that event A will occur is equal to 1 minus the probability that event A will not occur. P(A) = 1 - P(A')

OpenStudy (anonymous):

The probability that Event A or Event B occurs is equal to the probability that Event A occurs plus the probability that Event B occurs minus the probability that both Events A and B occur. P(A ∪ B) = P(A) + P(B) - P(A ∩ B))

OpenStudy (anonymous):

Invoking the fact that P(A ∩ B) = P( A )P( B | A ), the Addition Rule can also be expressed as P(A ∪ B) = P(A) + P(B) - P(A)P( B | A )

OpenStudy (anonymous):

this is the answer 0.16